OpenStudy (anonymous):

Hi! please explain how to get the answer : 1 divided by 1/5 thank you

OpenStudy (anonymous):

5

OpenStudy (anonymous):

1/(1/5)=5 or 1/(0.2) =5

OpenStudy (anonymous):

whole number divided by a fraction write the 1 as a fraction \[\frac{ 1 }{ 1 }\div \frac{ 1 }{ 5 }\] Then flip over the bottom and multiply it by the top \[\frac{ 1 }{ 1 } * \frac{ 5 }{ 1 }=\frac{ 5 }{ 1 }\] hence 5

OpenStudy (kc_kennylau):

Well you don't need to write the 1 as a fraction but well

OpenStudy (anonymous):

ikr